High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costingReq 1 Reg 2 Req ЗА Req 3B Compute the Predetermined Overhead Rates used in the Molding Department and the Painting DepartmentReq 1 Req 2 Req 3A Req 3B Compute the total overhead cost applied to Job 205. (Round Predetermined overhead rate to 2 decimReq 1 Req 2 Req ЗА Req 3B What would be the total manufacturing cost recorded for Job 205? (Round Predetermined overhead ratReq 1 Reg 2 Req ЗА Req 3B If the job contained 32 units, what would be the unit product cost? (Round Predetermined overhead
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 1

Molding = $ 5.10 per MH
Painting = $ 14.10 per DLHs

Working

Molding

Painting

A [given]

Total Fixed manufacturing Overhead cost

$232,200

$551,930

Allocation base

machine hours

direct labor hours

B [given]

Amount of allocation base

                        86,000

                                     56,900

C = A/B

Fixed Overhead rate

$                         2.70

$                                      9.70

D [given]

Variable overhead rate

$                         2.40

$                                      4.40

E = C+D

Predetermined Overhead rates

$                         5.10

$                                    14.10

  • Requirement 2

Total Overhead cost allocated = (310 MHs x $ 5.10) + (129 DLHS x $ 14.10)
= 1581 + 1819
= $ 3400

  • Requirement 3

Molding

Painting

Total

Direct Material Cost

$934

$1,220

$2,154

Direct Labor Cost

$700

$970

$1,670

Manufacturing Overhead cost applied

$1,581

$1,819

$3,400

Total manufacturing cost assigned

$3,215

$4,009

$7,224

  • Requirement 4
    Unit Product Cost = $ 7224 / 32 units
    = $ 225.75 per unit.